4) The number of hurricanes which hit the USA each year is Poisson with parameter 4. What is the probability that next year at least 3 hurricanes will hit the USA?
Answer #1

Probability Mass Function of Poisson Distribution with parameter lambda = 4 is give by:

P(X)=rac{e^{-4}4^{x}}{x!},

x = 0, 1, 2, ...

P(Xgeq3) = 1 - [p(X=0) + P(X=1) + P(X=2)]

-440 0.0183 ×1 -0.0183 P(X =0)--0! 1!

e-441 1! 0.0183 × 4 1! = 0.0733

e-442 P(X = 2)--21 0.0183 × 16 21 = 0.1465

So,

P(Xgeq3) = 1 - (0.0183 + 0.0733 + 0.1465) = 1 - 0.2381 = 0.7619

So,

Answer is:

0.7619